# MAESTRO: Multi-Agent Environment Shaping through Task and Reward Optimization

### v1

@@ Abstract
Cooperative Multi-Agent Reinforcement Learning (MARL) is fundamentally constrained by two design bottlenecks: the manual engineering of dense reward functions and the difficulty of constructing curricula that prevent local optima in high-dimensional state spaces. Existing solutions typically rely on static, hand-crafted heuristics that struggle to generalize. While Large Language Models (LLMs) have recently been applied to decision-making, most approaches utilize the “LLM-as-Agent” paradigm—placing the model directly in the control loop—which incurs prohibitive inference latency and token costs unsuitable for real-time systems.
In this work, we propose MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a framework that shifts the role of the LLM from real-time agent to high-level architect. MAESTRO treats the training process as a dual-loop optimization problem where the LLM dynamically designs both the task and the solving guidance. The system integrates two generative components: (i) a semantic curriculum generator that synthesizes diverse, context-rich scenarios based on agent performance, and (ii) an automated reward synthesizer that writes executable Python reward functions to guide agents through the evolving curriculum. These components optimize the training of a standard off-policy MARL backbone (MADDPG) without requiring LLM inference during execution.
@@
Our code is available at: https://github.com/BrianWu1010/MAESTRO.git.
@@ 1 Introduction
Cooperative Multi-Agent Reinforcement Learning (MARL) holds immense promise for solving complex decentralized control problems, from warehouse logistics to sensor network coordination. However, scaling MARL to realistic, non-stationary environments is persistently hindered by two fundamental design bottlenecks: curriculum design and reward engineering.
The first bottleneck, exploration, is often addressed by **Curriculum Learning (CL)**, which gradually increases task difficulty. However, constructing effective curricula requires deep domain expertise to ensure tasks are meaningfully diverse and effective [narvekar2020curriculum]. The second bottleneck concerns reward specification: the efficacy of MARL is brittle with respect to reward specification. Standard dense rewards are often misaligned with global objectives [zang2020metalight], while sparse global rewards fail to provide sufficient learning signal. Handcrafting reward functions that generalize across dynamic regimes remains a labor-intensive and error-prone process.
Large Language Models (LLMs) offer a potential remedy by encoding rich semantic priors that can automate these design processes. However, the prevailing integration paradigm—LLM-as-Agent—typically inserts the model directly into the control loop to select actions [lai2023llm, da2023llm, zhang2023trafficgpt]. While this yields strong zero-shot generalization, it incurs prohibitive inference latency and token costs, rendering it impractical for real-time control systems requiring sub-second decision making. Furthermore, this approach underutilizes the LLM’s capacity for high-level reasoning, relegating it to per-step tactical decisions.
Our Approach: MAESTRO.
In this work, we propose a paradigm shift from LLM-as-Agent to LLM-as-Architect. We introduce MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a generative meta-learning framework that uses the LLM offline to shape the training process of a conventional MARL backbone. We treat the training loop as a dual-optimization problem in which the LLM acts as a high-level designer. MAESTRO integrates two generative components: (i) a Semantic Curriculum Generator that synthesizes diverse, performance-driven task contexts, and (ii) an Automated Reward Synthesizer that produces executable Python reward functions dynamically tailored to the current curriculum stage.
By distilling semantic knowledge into executable training scaffolds (tasks and rewards), MAESTRO guides a standard MADDPG policy, enabling the final agents to retain the lightweight, low-latency characteristics required for deployment while benefiting from the broad reasoning capabilities of foundation models.
Figure 1: The MAESTRO framework workflow. The system operates as a dual-loop process: The LLM (Architect) designs the learning environment by generating curriculum contexts and synthesizing executable Python reward functions. The MARL agents (Learners) train within this shaped environment. Performance feedback triggers the Architect to adapt the curriculum and regenerate rewards, isolating the expensive LLM inference from the real-time execution loop.
@@
We validate MAESTRO on the challenging domain of Urban Traffic Signal Control (TSC), a prominent benchmark for MARL coordination [wei2019colight, zhang2019cityflow]. TSC specifically serves as a rigorous testbed for non-stationarity and coordination. Our contributions are:
  • A Generative Environment Shaping Framework. We formally propose MAESTRO, a method for contextual MARL that automates both task generation and reward synthesis, shifting the LLM’s role to an offline supervisor.
  • Controlled Ablation of Generative Components. We conduct a rigorous ablation comparing three configurations: a baseline LLM-driven curriculum with template rewards (A2), an alternative curriculum mode (A8), and the full MAESTRO system with LLM-generated reward shaping (A7). This design explicitly isolates the value of generative reward shaping from curriculum learning.
[…]
@@ 2.1 MARL for Decentralized Control
Multi-Agent Reinforcement Learning (MARL) is widely applied to cooperative control problems where agents share a global goal but make local decisions based on partial observations. TSC, a challenging instance of a Dec-POMDP, popularized the Centralized Training with Decentralized Execution (CTDE) paradigm [lowe2017maddpg], where agents learn a joint value function during training but act independently in deployment. Early applications of CTDE to traffic, such as CoLight [wei2019colight] and MPLight [chen2020mplight], focused on modeling inter-agent dependencies using mechanisms like graph attention and phase-pressure signals. These approaches assume a static, manually designed reward function and a fixed environment setup. Our work maintains the stability and real-time capability of the CTDE backbone (MADDPG) but allows the reward function itself and the training environment to evolve dynamically via LLM-generated shaping signals, addressing a critical generalization gap.
2.2 Automated Curriculum and Environment Generation
The difficulty of exploration in high-dimensional MARL motivates Automatic Curriculum Learning (ACL), which autonomously sequences tasks to maximize agent learning progress [portelas2020automatic]. Methods like POET [wang2019paired] and ALP-GMM [portelas2020automatic] demonstrate the power of generating diverse environments to force agents to acquire generalizable skills. In the context of TSC, curriculum generation has traditionally been limited to heuristic, numeric adjustments of traffic parameters, such as flow rates and phase durations [zang2020metalight]. These numeric methods lack the expressiveness to generate complex, semantically meaningful scenarios (e.g., “a peak-hour commuter pattern with a major bottleneck on the eastern artery”). Our approach moves beyond numeric scaling to a language-informed ACL, where an LLM generates semantically rich traffic contexts conditioned on a scalar difficulty signal, significantly expanding the scope of environment shaping.
2.3 LLMs for Reward and Policy Generation
The integration of Large Language Models into RL systems is broadly categorized into two paradigms:
LLM-as-Agent.
In this setting, the LLM directly selects actions or high-level plans, often via chain-of-thought reasoning. TrafficGPT [zhang2023trafficgpt] and LLM-TSCS [lai2023llm] exemplify this for traffic control. While providing strong zero-shot behavior, these methods are bottlenecked by the LLM’s high inference latency and token cost, making them ill-suited for real-time control applications with strict timing constraints.
@@ 2.3 LLMs for Reward and Policy Generation
This paradigm leverages the LLM’s reasoning capabilities during the training phase to shape the learning process. Recent work like Eureka [ma2023eureka] and Code-as-Policies [liang2023code] demonstrated that LLMs can generate effective policy code or, critically, executable reward functions for single-agent or robotic skill acquisition. These reward functions act as powerful, automatically generated priors.
MAESTRO builds upon and substantially extends the LLM-as-Teacher direction for the cooperative MARL domain. We introduce two key novelties: (i) the generative process produces executable Python reward functions for a decentralized, cooperative MARL system, which is significantly more complex than the single-agent setting; and (ii) we explicitly couple reward generation with a performance-driven, LLM-generated curriculum, ensuring the reward logic evolves in synergy with the increasing task difficulty. This coupling is essential for the stability and generalization of MARL policies in non-stationary environments.
@@ 3 Preliminaries
3.1 Contextual MARL Formulation and Objectives
We model multi-intersection TSC as a contextual Decentralized Partially Observable Markov Decision Process (Dec-POMDP), augmented with the MAESTRO training framework. The environment is defined as
@@ 3.1 Contextual MARL Formulation and Objectives
  • $\mathcal{C}$ denotes the curriculum context space (generated by the LLM Architect), which parameterizes traffic flows, route patterns, and demand distributions.
  • $\mathcal{I}=\{1,\dots,N\}$ is the set of traffic-signal agents (intersections).
@@
  • $\mathcal{R}_{C}$ is the context-dependent reward function (synthesized by the LLM Architect), and $\gamma$ is the discount factor.
At the start of each training stage, a context $C\in\mathcal{C}$ is sampled and the LLM Architect produces the corresponding reward function $\mathcal{R}_{C}$. Agents then interact with the environment under the joint policy $\pi_{C}$, generating a trajectory $\tau$. The objective is to maximize the expected cumulative return over all curriculum contexts:
$\max_{\pi}\;\mathbb{E}_{C\sim p(\mathcal{C}),\;\tau\sim\mathcal{T}_{C}(\cdot\mid\pi_{C})}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}_{C}(s_{t},a_{t})\right].$ (1)
We adopt MADDPG [lowe2017maddpg] as the MARL Learner backbone within the Centralized Training with Decentralized Execution (CTDE) paradigm. This design preserves stability during cooperative learning while maintaining efficient, decentralized execution at test time—cleanly separating the LLM’s offline design role from the agents’ real-time control.
3.2 MARL Learner Specification
@@
Experiments use the CityFlow microscopic simulator [zhang2019cityflow] on the real-world Hangzhou (HZ) network with $|\mathcal{I}|=16$ signalized intersections. Each episode spans 360 timesteps (1s per step). The curriculum context $C$ configures dynamic elements such as route distributions, vehicle arrival patterns, and vehicle parameters. Full parameter ranges are provided in the appendix.
@@
Each agent $i$ receives a structured observation vector $o_{i}^{t}$ comprising local queue lengths, inflow/outflow rates, current phase, elapsed phase duration, lane occupancy, and short-horizon traffic history. This feature design follows established MARL–TSC practice and is detailed in the appendix. Agent actions correspond to selecting one of the $|P_{i}|=4$ feasible signal phases at each timestep, consistent with intersection geometry.
@@
Performance is evaluated using standard TSC metrics: mean travel time, delay, queue length, and throughput. In MAESTRO, these metrics serve two purposes: (i) they drive difficulty adjustment for the curriculum generator, and (ii) they are combined into the composite reward signal $\mathcal{R}_{C}$ used during agent training.
@@ 4 Methodology
This section introduces MAESTRO and the three evaluated configurations. Figure 1 illustrates the overall workflow, and Figure 2 shows the conceptual foundation.
Figure 2: Conceptual foundation of the MAESTRO architecture. The LLM acts as the Architect, generating task-specific, executable components (e.g., reward functions and curriculum contexts) that are injected into the MARL training loop. MAESTRO focuses on LLM-generated reward shaping coupled with an adaptive curriculum, while keeping the underlying MARL Learner backbone (MADDPG) unchanged for deployment safety and real-time efficiency.
4.1 Overview of MAESTRO Configurations
We evaluate three configurations that differ only in how they utilize the LLM for generative curriculum design and reward shaping. All other components (network, MADDPG architecture, and hyperparameters) remain identical, enabling a clean controlled ablation:
@@ 4.1 Overview of MAESTRO Configurations
All configurations use LLM assistance to generate curriculum contexts, but only A7 (the full MAESTRO system) incorporates the LLM-generated auxiliary reward. The ablation focuses on isolating the contribution of generative reward shaping within an already LLM-assisted curriculum.
4.2 LLM-Driven Adaptive Curriculum

### v2

@@ Abstract
Cooperative multi-agent reinforcement learning (MARL) is constrained by two core design bottlenecks: dense reward engineering and curriculum design. Hand-crafted rewards often misalign with global objectives, while static curricula fail to prevent agents from getting trapped in local optima. Large language models (LLMs) have been used to address some of these issues, but most work adopts an LLM-as-Agent paradigm that keeps the model in the control loop, incurring prohibitive latency and token costs for real-time systems.
We propose MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), which shifts the LLM’s role from online controller to offline training architect. MAESTRO treats training as a dual-loop optimization problem: an LLM shapes the learning environment, while a standard MARL backbone optimizes the policy. The LLM instantiates two generative components: (i) a semantic curriculum generator that synthesizes diverse, performance-conditioned traffic scenarios, and (ii) an automated reward and policy synthesizer that fills executable Python templates to produce reward shaping signals and prior policy logits. Prior policies are used only as a regularizer during actor updates (via MSE penalties that decay over training); the deployed controller is a standalone MADDPG policy with no LLM calls. Reward shaping is additively combined with environment rewards, with a difficulty-dependent weight.
@@
Code is available at https://github.com/BrianWu1010/MAESTRO.git.
@@ 1 Introduction
Cooperative multi-agent reinforcement learning (MARL) is a natural framework for decentralized control problems such as warehouse logistics, sensor networks, and urban traffic signal control (TSC). In principle, centralized training with decentralized execution allows agents to coordinate via learned policies while acting on local observations. In practice, scaling MARL to realistic, non-stationary environments remains difficult due to two persistent design challenges: curriculum design and reward engineering.
Curriculum learning (CL) addresses exploration by exposing agents to gradually harder tasks. However, effective curricula require domain knowledge to ensure that tasks are both diverse and aligned with the target deployment regime [narvekar2020curriculum]. Reward design is equally critical: dense rewards may mis-specify long-term objectives [zang2020metalight], while sparse global rewards provide weak learning signals. Crafting rewards that generalize across dynamic regimes is labor-intensive and brittle.
LLMs offer rich semantic priors that could automate both curricula and rewards. Yet most existing integrations follow an LLM-as-Agent paradigm: the LLM directly selects actions or high-level decisions in the control loop [lai2023llm, da2023llm, zhang2023trafficgpt]. This can yield strong zero-shot behavior but incurs high latency and token costs, limiting its applicability to real-time, safety-critical systems. It also underuses the LLM’s strength as a high-level designer.
Our approach: LLM-as-Architect.
We propose MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), which repositions the LLM as an offline architect of the training process. Rather than emitting actions, the LLM designs the training environment: it generates curriculum contexts and fills parameterized templates for reward shaping and prior policy logits. A standard MADDPG backbone then optimizes a policy under this shaped environment and reward.
MAESTRO’s training loop has two coupled components:
  1. 1.
    A semantic curriculum generator that maps a scalar difficulty level and performance statistics into natural-language traffic descriptions, which are compiled into CityFlow configurations.
  2. 2.
    An automated reward and policy synthesizer that outputs parameters for pre-validated Python templates. These templates implement an auxiliary reward function and a prior policy that regularizes the actor via a time-decaying MSE penalty, without intervening in action selection.
The resulting policy remains lightweight and low-latency at deployment, while the LLM’s reasoning is fully exploited during training.
Figure 1: MAESTRO workflow. The LLM (Architect) generates curriculum contexts and parameterizes executable templates for reward shaping and prior policy logits. MARL agents (Learners) train under this shaped environment using MADDPG. Performance feedback drives curriculum updates and, in A7, periodic regeneration of reward and policy templates. LLM calls occur only during training, not deployment.
@@
We validate MAESTRO on urban TSC, a standard cooperative MARL benchmark [wei2019colight, zhang2019cityflow], and make three contributions:
  • Generative environment shaping. We formalize MAESTRO as a contextual MARL framework in which an LLM automates both task generation and reward synthesis while remaining entirely offline at deployment.
  • Controlled ablation of LLM components. We compare three configurations: A2 (adaptive LLM curriculum with template rewards), A8 (alternative LLM curriculum with template rewards), and A7 (full MAESTRO with adaptive curriculum and LLM-shaped rewards), isolating the effects of curriculum design and reward shaping.
  • Stability and risk-adjusted performance. Across four seeds, A7 achieves markedly higher stability than both A2 and A8 while still improving mean return. In domain metrics, MAESTRO configurations reduce delay by 20–28% and increase throughput by 6–9% relative to the baseline.
@@ 2.1 MARL for Decentralized Control
MARL is widely used for cooperative control with a shared global objective and local observations. Centralized training with decentralized execution (CTDE) [lowe2017maddpg, rashid2018qmix] mitigates non-stationarity during learning by training joint critics while deploying decentralized actors. Even under CTDE, state-of-the-art methods can be sensitive to configuration and hyperparameters [papoudakis2021benchmarking], motivating additional structural priors.
Urban TSC is a canonical MARL benchmark. CoLight [wei2019colight] and MPLight [chen2020mplight] use graph attention and pressure-based features to model inter-agent interactions and achieve strong performance on real networks. However, they assume fixed environment specifications and manually designed rewards. Reward shaping theory [ng1999policy] warns that naive shaping can induce suboptimal policies, and in practice TSC rewards remain heuristic. MAESTRO retains a CTDE backbone (MADDPG) but allows both curricula and rewards to evolve via LLM-generated signals.
2.2 Automatic Curriculum and Environment Generation
Curriculum learning sequences tasks by difficulty to accelerate training [bengio2009curriculum]. Automatic curriculum learning (ACL) extends this idea by adapting task difficulty based on learning progress [portelas2020automatic]. Adversarial and co-evolutionary approaches such as PAIRED [dennis2020emergent] and POET [wang2019paired] construct challenging environments that drive robust skill acquisition.
In cooperative MARL, difficulty-aware curricula (e.g., cMALC-D [kim2023cmalcd]) adjust task parameters based on performance. In TSC, curricula typically perturb numeric parameters such as arrival rates or demand scales [zang2020metalight], which improves learning but captures only a narrow slice of real-world structure (e.g., complex rush-hour patterns or localized bottlenecks).
MAESTRO extends ACL in two ways. First, it leverages an LLM to generate semantically rich traffic descriptions, which are compiled into CityFlow configs, expanding the curriculum space beyond low-dimensional numeric perturbations. Second, reward shaping and curriculum progression are coupled: as difficulty changes, the LLM simultaneously revises environment descriptions and reward logic, aligning training signals with the evolving task distribution.
2.3 LLMs for Reward and Training-Process Generation
Two main paradigms integrate LLMs with RL.
LLM-as-Agent (online control).
Here, the LLM selects actions or high-level plans at inference time. TrafficGPT [zhang2023trafficgpt] and LLM-TSCS [lai2023llm] apply this to traffic management; LA-Light [lai2024lalight], Voyager [wang2023voyager], and Code-as-Policies [liang2023code] do so in robotics and embodied environments. These systems exploit LLM priors for zero-shot generalization but inherit high latency, token costs, and limited action frequency, making them problematic for real-time multi-intersection TSC.
@@ 2.3 LLMs for Reward and Training-Process Generation
An alternative is to use LLMs to shape training rather than deployment. Eureka [ma2023eureka] prompts an LLM to generate and refine reward functions for robotic tasks, improving policies without using the LLM at test time. In traffic control, Da Lio et al. [da2023llm] inject semantic priors via language, and LAMARL [lin2024lamarl] uses language-based heuristics to guide MARL with curriculum hints and shaping.
Within this teacher regime, an important distinction is whether the LLM remains in the deployed controller. MAESTRO belongs to the class where the runtime agent is purely an RL policy: the LLM is used only during training to generate curricula, reward functions, and policy priors. Compared to prior work, MAESTRO (i) applies template-based LLM reward synthesis to a cooperative 16-agent TSC setting, and (ii) tightly couples LLM-shaped rewards with an adaptive, performance-driven curriculum. We show that this joint environment–reward generation is key to the stability and risk-adjusted gains observed.
@@ 3 Preliminaries
3.1 Contextual MARL Formulation
We model multi-intersection TSC as a contextual Decentralized Partially Observable Markov Decision Process (Dec-POMDP), augmented with curriculum contexts generated by the LLM Architect:
@@ 3.1 Contextual MARL Formulation
  • $\mathcal{C}$ is the curriculum context space (LLM-generated traffic patterns).
  • $\mathcal{I}=\{1,\dots,N\}$ indexes intersections.
@@
  • $\mathcal{R}_{C}$ is the context-dependent reward function; $\gamma$ is the discount factor.
At each training stage, a context $C\in\mathcal{C}$ is sampled, and the LLM synthesizes $\mathcal{R}_{C}$. Agents then follow a joint policy $\pi_{C}$ to generate a trajectory $\tau$. The objective is to maximize expected return over the curriculum distribution:
$\max_{\pi}\;\mathbb{E}_{C\sim p(\mathcal{C}),\tau\sim\mathcal{T}_{C}(\cdot|\pi_{C})}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}_{C}(s_{t},a_{t})\right].$ (1)
We adopt MADDPG [lowe2017maddpg] as the MARL learner under CTDE, ensuring stable cooperative learning and fast decentralized execution.
3.2 Environment and Agent Specification
@@
We use CityFlow [zhang2019cityflow] on the real-world Hangzhou (HZ) network with 16 signalized intersections. Each episode spans 360 steps (1-second simulation interval). Context $C$ controls route distributions, arrival rates, and vehicle parameters; ranges are detailed in the appendix.
@@
Each agent $i$ observes a vector $o_{i}^{t}$ including local queue lengths, inflow/outflow, active phase, elapsed phase duration, occupancy, and short-term history, following standard MARL-TSC practice. Actions select one of four feasible signal phases per intersection.
@@
We track mean travel time, delay, queue length, and throughput. These metrics both guide curriculum difficulty updates and form components of the composite reward $\mathcal{R}_{C}$ used for training.
@@ 4 Methodology
Figure 1 illustrates MAESTRO’s dual-loop training. The outer loop updates curriculum difficulty and (in A7) triggers LLM regeneration of reward and policy templates; the inner loop runs standard MADDPG updates under the current shaped reward.
Figure 2: Conceptual foundation of MAESTRO. The LLM (Architect) generates task-specific components—curriculum contexts, reward parameters, and prior policy parameters—that are injected into the MARL training loop. The deployed controller is an unmodified MADDPG policy; LLM inference is restricted to training.
4.1 Configurations
We evaluate three configurations that share the same network architecture, hyperparameters, and simulator, differing only in their use of the LLM:
@@ 4.1 Configurations
All settings use LLM-generated curricula; only A7 adds LLM-shaped reward and policy priors.
Curriculum modes.
  • llm_adaptive (A2, A7): Structured prompts with explicit difficulty mapping and conservative regeneration triggers; parameters are clipped to bounded ranges.
  • llm (A8): Less constrained prompts allowing more diverse contexts, increasing expressiveness but also variability.
Both modes update a scalar difficulty $d\in[0.3,1.0]$, but llm tends to produce more volatile curriculum trajectories.
4.2 Adaptive Curriculum